OpenStudy (anonymous):

The perimeter of a triangle is 42 centimeters. The longest side is 2 centimeters less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of each side of triangle please

OpenStudy (anonymous):

let x denote the length of the longest side and y and z the othe rtwo sides we have y+z -2 = x and perimeter = x+y+z = 42 that is x + x+2=42 that is x = 20 and 2y = x-4 that is 2y = 16 that is y = 8 and z = 14